The ledger of Ruru Company at the end of the current year shows Accounts Receivable $200,000, Sales $1,400,000, and Sales Returns and Allowances $50,000. Instructions:

(a) If Ruru uses the direct write-off method to account for uncollectible accounts, journalize theadjusting entry at December 31, assuming Ruru determines that Barking Ghosts Company’s $2,400 balance is uncollectible.

(b) If Allowance for Doubtful Accounts has a credit balance of $3,500 in the trial balance, journalize the adjusting entry at December 31, assuming bad debts are expected to be (1) 1% of net sales, and (2) 10% of accounts receivable.

(c) If Allowance for Doubtful Accounts has a debit balance of $370 in the trial balance, journalize the adjusting entry at December 31, assuming bad debts are expected to be (1) 0.75% of net sales and (2) 6% of accounts receivable.

(d) Why is the allowance method required by GAAP?
Answer #1
Notes 1
Allowance of doubtful account (credit balance) after adjustment (200000*10%)      20,000

Less: Allowance of doubtful account (credit balance) available before adjustment

(As per question) (credit balance means already adjustment made.)

     (3,500)
Adjustment of bad debts expense      16,500
Notes 2
Allowance of doubtful account (credit balance) after adjustment (200000*6%)      12,000

Add: Allowance of doubtful account (debit balance) available before adjustment

(As per question) (debit balance means extra adjustment required to make.)

           370
Adjustment of bad debts expense      12,370
If Ruru uses the direct write-off method to account for uncollectible accounts, journalize the adjusting entry at December 31, assuming Ruru determines that Barking Ghosts Company’s $2,400 balance is uncollectible.
Adjustment entries
Transaction General journal Debit credit
a Bad debts expense          2,400
Accounts receivable - Barking Ghosts Company          2,400
To record bad debts expense.
If Allowance for Doubtful Accounts has a credit balance of $3,500 in the trial balance, journalize the adjusting entry at December 31, assuming bad debts are expected to be (1) 1% of net sales, and (2) 10% of accounts receivable.
Adjustment entries
Transaction General journal Debit credit
b1 Bad debts expense       13,500
Allowance of doubtful account       13,500
To record Bed debts expense under Sales method. (net sales = 1400000-50000 = 1350000) (1350000*1%)
b2 Bad debts expense       16,500
Allowance of doubtful account       16,500
To record Bed debts expense. (notes 1)
If Allowance for Doubtful Accounts has a debit balance of $370 in the trial balance, journalize the adjusting entry at December 31, assuming bad debts are expected to be (1) 0.75% of net sales and (2) 6% of accounts receivable.
Adjustment entries
Transaction General journal Debit credit
c1 Bad debts expense       10,125
Allowance of doubtful account       10,125
To record Bed debts expense under Sales method. (net sales = 1400000-50000 = 1350000) (1350000*0.75%)
c2 Bad debts expense       12,370
Allowance of doubtful account       12,370
To record Bed debts expense. (notes 2)
Why is the allowance method required by GAAP?
As per the matching concept, revenues and expenses related to revenue should record in the same accounting period. The allowance method helps to confirm the U.S.GAAP by matching concept through estimating bad debts expenses and recorded in the same accounting period.